OpenStudy (anonymous):

I need help with some factoring problems. I want to check my work. 1. u^3=14u^2+32u 2. 2d^3+6d=7d^2 (The ^ is an exponent)

OpenStudy (campbell_st):

well what did you get for both questions...?

OpenStudy (anonymous):

\[u ^{3}-14u ^{2}-32u=u \left( ? \right)\]

OpenStudy (anonymous):

for #1. I got -u(u^2-14u-32)=0 which factors to -u(u+16)(u-2)=0

OpenStudy (anonymous):

for #2 I can't get past d(2d^2-7d+6)=0

OpenStudy (anonymous):

The factors are d, 2d-3, and d-2.

OpenStudy (anonymous):